# Extending free actions of finite groups on unoriented surfaces

### v4

@@ Introduction
The Schur and Bogomolov multipliers have been of great interest because of the solution in [ASSU23] of the following question: Are all free actions on oriented surfaces equivariantly bound? There is an affirmative answer in [RZ96, DS22] for abelian, dihedral, symmetric, and alternating groups. However, we can find counterexamples in [Sam22] for finite groups with non-trivial Bogomolov multiplier. For a finite group $G$, the Bogomolov multiplier $B_{0}(G)$ is trivial when toral classes generate the Schur multiplier. The extension to a 3-manifold has a set of fixed points, called the ramification locus, given by a link of circles. It was believed that if $B_{0}(G)\neq 0$, then the extension can be the gluing of spherical actions by dihedral groups $D_{2n}$, the alternating groups $A_{4}$, $A_{5}$, or the symmetric group $S_{4}$. The ramification locus is called singular if it fails to be a submanifold of the 3-manifold, i.e., an embedded graph of fixed points. Notwithstanding, it was checked in [ASSU23] that applying the blow-up construction can replace the singular ramification locus with a non-singular ramification locus, which has only cyclic or dihedral groups as isotropy groups. Additionally, the points with dihedral isotropy type are cancelled by attaching certain equivariant handles, described in [ASSU23, Sec. 4]. In conclusion, the Bogomolov multiplier is a complete obstruction to extending finite free actions from closed, oriented surfaces to 3-manifolds.
@@
The unoriented case affords another type of relation since we have to consider the set of squares $S(G)=\langle x^{2}:x\in G\rangle$, which are the associated monodromies for the Möbius band. We denote by $\bigl(G,G\bigl)$ the free group on pairs $\langle x,y\rangle$, $(x^{\prime},y^{\prime})$ and $(z)$, with $x,y,x^{\prime},y^{\prime},z\in G$. The elements $\langle x,y\rangle$ are the oriented pairs sent in $S(G)$ to the commutator $[x,y]$. The elements $(x,y)$ are called the unoriented pairs that are sent in $S(G)$ to the unoriented commutator $\{x,y\}:=xyx^{-1}y$. The elements $(z)$ are the square elements sent in $S(G)$ to $z^{2}$. Consequently, we obtain a canonical map $\bigl(G,G\bigl)\longrightarrow S(G)$, defined by $\langle x,y\rangle\longmapsto[x,y]$, $(x,y)\longmapsto xyx^{-1}y$, and $(z)\longmapsto z^{2}$. The unoriented Schur multiplier is the quotient
$\mathcal{M}(G;\mathbb{Z}_{2}):=\frac{\operatorname{ker}\left(\bigl(G,G\bigl)\longrightarrow S(G)\right)}{N}\,$,
where $N$ is the normal subgroup generated by the following seven relations:
@@
$(y^{x},z^{x})^{-1}$ $\sim(x,\{y,z\}^{-1})(y,z)\,.$
We can show that the oriented relations in Theorem 2 are a consequence of the unoriented relations; see Theorem 8. However, we do not know if we can reduce the number of unoriented relations to fewer than seven. We obtain a Hopf formula as follows:
@@
Unlike the oriented case, we now have more extendable elements. In addition to the toral classes, we have free actions with the quotient of a Klein surface, which is extendable by the same reasons as the torus. However, there is an extra extendable element. Consider the principal $G$-bundles over the projective space $\mathbb{R}P^{2}$, which are of two types: the trivial bundle one which we will consider by default a trivial element, and those which correspond to non-trivial 2-torsion elements in the group, i.e., the elements $x\in G$ with $x^{2}=1$ and $x\neq 1$. These elements are extendable because the total space is the disjoint union of $|G|/2$ spheres. Dropping the subgroup generated by the trivial $G$-bundle over $\mathbb{R}P^{2}$, we define the unoriented Bogomolov multiplier, which we denote by $B_{0}(G;\mathbb{Z}_{2})$, as the quotient of the unoriented Schur multiplier $\mathcal{M}(G;\mathbb{Z}_{2})$ by the subgroup generated by free actions where the quotient has the possibilities to be a torus, a Klein bottle, and a projective space $\mathbb{R}P^{2}$.
We show that the unoriented Bogomolov multiplier is trivial for finite abelian, dihedral, symmetric, and alternating groups. The most important contribution of the present paper is an example of a group where the unoriented Bogomolov multiplier is not trivial: this is a group of order $64$ which is the fist group with non-trivial Bogomolov multiplier; see Example 4.1. Therefore, this is an example of a free surface action that does not equivariantly bounds even considering non-necessarily free actions over unoriented 3-manifolds.
@@ Acknowledgment
We would like to thank Eric Samperton and Bernardo Uribe for relevant conversations regarding this work. We also thank the financial support of “Convocatoria PAEP Posgrado de Matemáticas de la UNAM 2023.” The third author is supported by Investigadores por México SECIHTI.
@@ 1 $G$-cobordism
A $G$-cobordism over a Klein bottle depends on two elements $x,y\in G$, with boundary monodromy ${x,y}:=xyx^{-1}y$. Two equivalent representations are shown in Figure 5.
@@
Examples of extendable $G$-cobordisms include trivial bundles over orientable surfaces, which extend to solid handlebodies. A non-example is the trivial bundle over $\mathbb{R}P^{2}$. The following proposition generalizes extension constructions to certain unoriented surfaces.
@@ 2 The unoriented Schur multiplier
Isaai Schur [Sch04] introduced a mathematical object to classify projective representations $\rho\colon G\longrightarrow\operatorname{Gl}(n,\mathbb{C})/Z$, where $Z$ denotes the center of scalar matrices, using the quotient of $2$-cocycles by the group of $2$-coboundaries. Initially termed the Multiplikator of $G$, it is now known as the Schur multiplier $\mathcal{M}(G)$. We refer to [Wi81] for a historical overview and discussion. Over time, this object acquired interpretations in terms of group homology and cohomology [Br82], and it later became significant in low-dimensional topology, where it classifies free actions of finite groups on closed oriented surfaces (see [Ed82, Ed83]). In summary, $\mathcal{M}(G)$ can be identified with three equivalent groups: the second homology group $H_{2}(G,\mathbb{Z})$, the second cohomology group $H^{2}(G,\mathbb{C}^{*})$, and the second free oriented bordism group $\Omega_{2}^{SO}(G)$.
@@
Let $K^{\prime}:=\ker\bigl(\bigl(G,G\bigr)\to S(G)\bigr)$ be the kernel of the canonical map. The unoriented Schur multiplier, denoted $\mathcal{M}(G;\mathbb{Z}_{2})$, is defined as the quotient of $K^{\prime}$ by the normal subgroup generated by the following seven relations:
@@
$(y^{x},z^{x})^{-1}$ $\sim(x,\{y,z\}^{-1})(y,z)\,.$ (18)
@@
The substitution of $a=xy$, $b=y^{-1}$ implies the identity. Finally, (27) is obtained from the right-hand side of (26). ∎
We also encounter the relation $(z)\langle x,y\rangle\sim\langle y,x\rangle^{z}(z[x,y])$, which appears in [KT17] in the context of equivariant Klein TQFT, but whose derivation from the defining relations remains open.
@@
As in the oriented case, the argument reduces to groups with finitely many generators. Proposition 9 handles the one-generator case. Now let G=ABG=A*B with $|A|,|B|<\infty$, and assume $\mathcal{M}(A;\mathbb{Z}_{2})=1$ and $\mathcal{M}(B;\mathbb{Z}_{2})=1$. By induction, we may take $B$ to be free on a single generator $b$.
Using (13) and (14), every symbol can be expressed in terms of square elements. Moreover, the identity
$(z_{1}z_{2})\sim\langle z_{1}^{-1},z_{1}z_{2}\rangle^{z_{1}}(z_{1})(z_{2})\,$,
together with results on the oriented Schur multiplier [Kar87, Lemma 2.6.2, pp. 68-69] and repeated application of (15), shows that any element of $(G,G)$ can be written as a product of factors of the form $(a)$, $(a^{\prime}b)$, and $(b)$ with $a,a^{\prime}\in A$, $b\in B$.
Let $\omega\in(G,G)$ be a reduced word in these symbols, using (12) to combine powers of the same generator. By induction on the length of $\omega$, one checks that $\omega$ corresponds to a reduced word in the free product ABA*B, with possible endings as listed below:
Last two symbols in $\omega$ Corresponding word in $G$
$(b)(a)$ $b^{2}a^{2}$
$(ab)(a)$ $ababa^{2}$
$(a_{1})(a_{2})$, with $a_{1}\neq a_{2}$ $a_{1}^{2}a_{2}^{2}$, $a_{1}\neq a_{2}$
$(a)(b)$ $a^{2}b^{2}$
$(ab)(b)$ $abab^{3}$
$(a_{1}b)(a_{2}b)$ $a_{1}ba_{1}ba_{2}ba_{2}b$
$(b)(ab)$ $b^{2}abab$
$(a_{1})(a_{2}b)$ $a_{1}^{2}a_{2}ba_{2}b$
Words of the third type could continue to the left without the letter $b$, but inductively, we can exclude this possibility. Consequently, $\omega]\neq 1$ whenever $\omega$ is nonempty, proving the theorem. ∎
@@ 3 The Hopf formula for the unoriented case
Recall that the unoriented Schur multiplier is defined as the quotient of the kernel $K^{\prime}:=\ker\bigl((G,G)\to S(G)\bigr)$ by the normal subgroup $N$ generated by relations (12)–(18). Given a square-central extension $1\rightarrow A\rightarrow E\stackrel{{\scriptstyle\eta}}{{\rightarrow}}G\rightarrow 1$ (see Definition 13), we define a homomorphism $(G,G)\to E$ by mapping the generators $\langle x,y\rangle$, $(x^{\prime},y^{\prime})$, and $(z)$ to $\overline{x},\overline{y}]$, $\overline{x}^{\prime},\overline{y}^{\prime}$}, and $\overline{z}^{2}$, respectively, where $\eta(\overline{x})=x$, $\eta(\overline{y})=y$, $\eta(\overline{x}^{\prime})=x^{\prime}$, $\eta(\overline{y}^{\prime})=y^{\prime}$, and $\eta(\overline{z})=z$. Because the extension is square-central, this homomorphism is well defined. Moreover, its restriction to $K^{\prime}$ has image contained in $S(E)\cap A$, and it sends $N$ to $1$, inducing a surjective homomorphism $\mathcal{M}(G;\mathbb{Z}_{2})\to S(E)\cap A$. Consequently, we obtain the exact sequence
@@ 4 The unoriented Bogomolov multiplier
For a finite group $G$, the unoriented Bogomolov multiplier, denoted $B_{0}(G;\mathbb{Z}_{2})$, is the quotient of the unoriented Schur multiplier $\mathcal{M}(G;\mathbb{Z}_{2})$ by the subgroup generated by the $G$-cobordisms over the torus $\langle x,y\rangle$, the Klein bottle $(x^{\prime},y^{\prime})$, and the projective space $(z)$, subject to $[x,y]=1$, ${x^{\prime},y^{\prime}}=1$, and $z^{2}=1$.
@@
The known examples of groups with nontrivial Bogomolov multiplier $B_{0}(G)$ and even order satisfy $|G|=2^{n}$. The first such examples occur at order $2^{6}$; in GAP [GAP] they correspond to $\mathsf{SmallGroup}(64,i)$ for $i\in{149,150,151,170,171,172,177,178,182}$. Additionally, there are precisely eleven isoclinism families of groups of order $128$ whose Bogomolov multipliers are nontrivial. A distinctive feature of these examples is that $B_{0}(G)$ is always an elementary abelian $2$-group. Therefore, if $B_{0}(G)$ is an elementary abelian $2$-group for a group $G$ of order $2^{n}$, we obtain $B_{0}(G;\mathbb{Z}_{2})\cong B_{0}(G)$. This property is confirmed for extraspecial $2$-groups of order $2^{2n+1}$ with $n>1$, where the Schur multiplier is an elementary abelian $2$-group of order $2^{2n^{2}-n-1}$; see [Kar87, p. 138]. A potential counterexample could be a group of order $2^{n}$ that provides a counterexample to Schur’s conjecture [BKW73] (Urban Jezernik, personal communication).
@@
For finite abelian, dihedral, symmetric groups, and alteranting groups the unoriented Bogomolov multiplier is trivial.
@@
Using [GAP], one can verify that the first integral homology are the image of two subgroups isomorphic $\mathbb{Z}_{4}$ and one $\mathbb{Z}_{8}$. Moreover, the image of the second integral homology of every proper subgroup is trivial. Consequently, the class (39) remains nontrivial in the unoriented Bogomolov multiplier $B_{0}(\mathbb{Z}_{8}\rtimes Q_{8};\mathbb{Z}_{2})$. The reader may observe that, by applying the isomorphism (38) and that $B_{0}(G)$ is an elementary abelian 2-group, one immediately concludes that the unoriented Bogomolov multiplier is nontrivial.

### v5

@@ Introduction
The Schur and Bogomolov multipliers have been of great interest because of the solution in [ASSU23] of the following question: Do all free actions on oriented surfaces equivariantly bound? There is an affirmative answer in [RZ96, DS22] for abelian, dihedral, symmetric, and alternating groups. However, we can find counterexamples in [Sam22] for finite groups with non-trivial Bogomolov multiplier. It was shown in [ASSU23] that a finite free action on a closed, oriented surface extends to a (not necessarily free) action on a $3$-manifold if and only if the corresponding class in the Bogomolov multiplier $B_{0}(G)$ is trivial.
@@
The unoriented case affords another type of relations since we have to consider the set of squares $S(G)=\langle x^{2}:x\in G\rangle$, which are the associated monodromies for the Möbius band. We denote by $\bigl(G,G\bigr)$ the free group on pairs $\langle x,y\rangle$, $(x^{\prime},y^{\prime})$ and $(z)$, with $x,y,x^{\prime},y^{\prime},z\in G$. The elements $\langle x,y\rangle$ are the oriented pairs sent in $S(G)$ to the commutator $[x,y]$. The elements $(x,y)$ are called the unoriented pairs that are sent in $S(G)$ to the unoriented commutator $\{x,y\}:=xyx^{-1}y$. The elements $(z)$ are the square elements sent in $S(G)$ to $z^{2}$. Consequently, we obtain a canonical map $\bigl(G,G\bigl)\longrightarrow S(G)$, defined by $\langle x,y\rangle\longmapsto[x,y]$, $(x,y)\longmapsto xyx^{-1}y$, and $(z)\longmapsto z^{2}$. The unoriented Schur multiplier is the quotient
$\mathcal{M}(G;\mathbb{Z}_{2}):=\frac{\operatorname{ker}\left(\bigl(G,G\bigl)\longrightarrow S(G)\right)}{N^{\prime}}\,$,
where $N^{\prime}$ is the normal subgroup generated by the following relations:
@@
$(y^{x},z^{x})^{-1}$ $\sim(x,\{y,z\}^{-1})(y,z)\,$,
$(z)\langle x,y\rangle$ $\sim\langle y,x\rangle^{z}(z[x,y])\,.$
and two commutation relations:
$(b)\langle a_{0},b_{0}\rangle\sim\langle a_{0},b_{0}\rangle^{b^{2}}(b)\,$,
$\langle a_{0},b_{0}\rangle(a)\sim(a)\langle a_{0},b_{0}\rangle^{a^{-2}}\,.$
We can show that the oriented relations in Theorem 2 are a consequence of the unoriented relations; see Theorem 8. We obtain a Hopf formula as follows:
@@
The Bogomolov multiplier is defined as the quotient of the Schur multiplier $\mathcal{M}(G)$ by the toral classes provided by pairs of commuting elements. These toral classes are always extendable elements, as shown in [DS22, Sam20]. Unlike the oriented case, we now have more extendable elements; see Proposition 4. More precisely, in addition to the toral classes, we have free actions whose quotient is a Klein surface, which is extendable for the same reasons as the torus, as well as principal $G$-bundles over the projective space $\mathbb{R}P^{2}$ induced by elements $x\in G$ with $x^{2}=1$ and $x\neq 1$. By dropping the subgroup generated by the trivial $G$-bundle over $\mathbb{R}P^{2}$, we define the unoriented Bogomolov multiplier, which we denote by $B_{0}(G;\mathbb{Z}_{2})$, as the quotient of the unoriented Schur multiplier $\mathcal{M}(G;\mathbb{Z}_{2})$ by the subgroup generated by free actions where the quotient can be a torus, a Klein bottle, or a projective space $\mathbb{R}P^{2}$.
We show that the unoriented Bogomolov multiplier is trivial for finite abelian, dihedral, symmetric, and alternating groups. The most important contribution of the present paper is an example of a group where the unoriented Bogomolov multiplier is not trivial: this is a group of order $64$ which is the first group with non-trivial Bogomolov multiplier; see Example 4.1. In the forthcoming work [Seg], we will prove Theorem 19, which therefore implies that such a group acts in a freely way that does not equivariantly bound, even if we allow for non-free and unoriented preserving actions.
@@ Acknowledgment
We would like to thank Eric Samperton and Bernardo Uribe for relevant conversations regarding this work. We also thank the financial support of “Convocatoria PAEP Posgrado de Matemáticas de la UNAM 2023.” The third author is supported by Investigadores por México SECIHTI. We thank the anonymous reviewer for the careful reading and valuable suggestions that helped improve the manuscript.
@@ 1 $G$-cobordism
A $G$-cobordism over a Klein bottle depends on two elements $x,y\in G$, with boundary monodromy $\{x,y\}:=xyx^{-1}y$. Two equivalent representations are shown in Figure 5.
@@
Examples of extendable $G$-cobordisms include trivial bundles over orientable surfaces, which extend to solid handlebodies. A non-example is the trivial bundle over $\mathbb{R}P^{2}$ when $G$ is of odd order. The following proposition generalizes extension constructions to certain unoriented surfaces.
@@ 2 The unoriented Schur multiplier
Issai Schur [Sch04] introduced a mathematical object to classify projective representations $\rho\colon G\longrightarrow\operatorname{Gl}(n,\mathbb{C})/Z$, where $Z$ denotes the center of scalar matrices, using the quotient of $2$-cocycles by the group of $2$-coboundaries. Initially termed the Multiplikator of $G$, it is now known as the Schur multiplier $\mathcal{M}(G)$. We refer to [Wi81] for a historical overview and discussion. Over time, this object acquired interpretations in terms of group homology and cohomology [Br82], and it later became significant in low-dimensional topology, where it classifies free actions of finite groups on closed oriented surfaces (see [Ed82, Ed83]). In summary, $\mathcal{M}(G)$ can be identified with three equivalent groups: the second homology group $H_{2}(G,\mathbb{Z})$, the second cohomology group $H^{2}(G,\mathbb{C}^{*})$, and the second free oriented bordism group $\Omega_{2}^{SO}(G)$.
@@
Let $K^{\prime}:=\ker\bigl(\bigl(G,G\bigr)\to S(G)\bigr)$ be the kernel of the canonical map. The unoriented Schur multiplier, denoted $\mathcal{M}(G;\mathbb{Z}_{2})$, is defined as the quotient of $K^{\prime}$ by the normal subgroup generated by the following relations:
@@
$(y^{x},z^{x})^{-1}$ $\sim(x,\{y,z\}^{-1})(y,z)\,$, (18)
$(z)\langle x,y\rangle$ $\sim\langle y,x\rangle^{z}(z[x,y])\,.$ (19)
The relation (19) appears in [KT17] in the context of equivariant Klein TQFT. In addition, we have two commutation relations:
$(b)\langle a_{0},b_{0}\rangle\sim\langle a_{0},b_{0}\rangle^{b^{2}}(b)\,$, (20)
$\langle a_{0},b_{0}\rangle(a)\sim(a)\langle a_{0},b_{0}\rangle^{a^{-2}}\,.$ (21)
These commutation relations correspond to two equivalent ways of describing the $G$-cobordism over the pair of pants with inputs $x$ and $y$; see Figure 8 for an illustration.
Figure 8: Two equivalent ways of describing the $G$-cobordism over the pair of pants with inputs $x$ and $y$ and output xyxy.
@@
The substitution of $a=xy$, $b=y^{-1}$ implies the identity. Finally, the left-hand side from (30) is obtained from the right-hand side of (29) and right-hand side from (30) is obtained from relation (19). ∎
@@
As in the oriented case, the argument reduces to groups with finitely many generators. Proposition 9 handles the one-generator case. Now let G=ABG=A*B with $|A|,|B|<\infty$, and assume $\mathcal{M}(A;\mathbb{Z}_{2})=1$ and $\mathcal{M}(B;\mathbb{Z}_{2})=1$.
Firstly, using relations (13) and (14) together with (26) and (27), we obtain that any element of $(G,G)$ can be written as a product of factors of the form $(a)$, $(a^{\prime}b)$, and $(b)$, with $a,a^{\prime}\in A$, $b\in B$. Nevertheless, using (26) and (27) again, we obtain that every element in $(G,G)$ is the product of elements of $(a)$, $(b)$, and $\mu$, where $\mu$ lies in the subgroup of $(G,G)$ generated by all $\langle a,b\rangle$ with $a\neq 1$ in $A$ and $b\neq 1$ in $B$.
Similarly to the arguments in [Mi52] (or see [Kar87, Lemma 2.6.2, pp. 68-69]), we need the commutation identities similar to (8), (9), and (10), together with repeated use of the rules:
$\langle a,a^{\prime}\rangle^{b_{0}}\sim\langle b_{0},[a,a^{\prime}]\rangle\langle a,a^{\prime}\rangle$ (31)
$\langle a,b\rangle^{a_{0}}\sim\langle a_{0}a,b\rangle\langle b,a_{0}\rangle$ (32)
$\langle a,b\rangle^{b_{0}}\sim\langle b_{0},a\rangle\langle a,b_{0}b\rangle$ (33)
and three similar rules, obtained from these by interchanging $a$ with $b$, $a_{0}$ with $b_{0}$, and $a^{\prime}$ with $b^{\prime}$ (see [Kar87, p. 69]).
Recall the commutation relations (20) and (21) given by $(b)\langle a_{0},b_{0}\rangle\sim\langle a_{0},b_{0}\rangle^{b^{2}}(b)$ and $\langle a_{0},b_{0}\rangle(a)\sim(a)\langle a_{0},b_{0}\rangle^{a^{-2}}$. For the commutation between elements $(a)$ and $(b)$, we use:
$(a)(b)(a^{-1})(b^{-1})$ $\sim\langle a,ab\rangle(ab)(a^{-1})(b^{-1})$ (34)
$\sim\langle a,ab\rangle\langle ab,aba^{-1}\rangle(aba^{-1})(b^{-1})$ (35)
$\sim\langle a,ab\rangle\langle ab,aba^{-1}\rangle\langle aba^{-1},[a,b]\rangle([a,b])$ (36)
$\sim\langle a,ab\rangle\langle ab,aba^{-1}\rangle\langle aba^{-1},[a,b]\rangle\langle a,b\rangle^{2}$ (37)
$\sim\langle a,b\rangle^{a}\langle b,a^{-1}\rangle^{aba}\langle a^{-1},b^{-1}\rangle^{ab^{2}}\langle a,b^{-1}\rangle^{aba^{-1}}\langle a,b\rangle^{2}$ (38)
where we used the right-hand side of (30) in relation (37). Consequently, by applying relations (31), (32), (33), and the other three, we obtain the last commutation:
$(b)(a)\sim\mu^{\prime}(a)(b)$ (39)
where $\mu^{\prime}$ belongs to the subgroup of $(G,G)$ generated by all $\langle a,b\rangle$ with $a\neq 1$ in $A$ and $b\neq 1$ in $B$.
Using the commutation relations (20), (21), and (39), together with the simplification of the conjugations (31), (32), (33), and the other three, we arrive at the conclusion that every word $\omega\in(G,G)$ can be reduced to $\omega\sim\alpha\mu\beta$, where $\alpha$ is a word composed of elements $(a)$, $\langle a^{\prime},a^{\prime\prime}\rangle$, with $a,a^{\prime},a^{\prime\prime}\neq 1$; $\beta$ is a word composed of elements $(b)$, $\langle b^{\prime},b^{\prime\prime}\rangle$, with $b,b^{\prime},b^{\prime\prime}\neq 1$; and $\mu$ lies in the subgroup of $(G,G)$ generated by all $\langle a,b\rangle$ with $a\neq 1$ in $A$ and $b\neq 1$ in $B$. As in Miller’s proof, we obtain $\omega]\neq 1$ whenever $\omega$ is nonempty, proving the theorem. ∎
@@ 3 The Hopf formula for the unoriented case
Recall that the unoriented Schur multiplier is defined as the quotient of the kernel $K^{\prime}:=\ker\bigl((G,G)\to S(G)\bigr)$ by the normal subgroup $N^{\prime}$ generated by relations (12)–(18). Given a square-central extension $1\rightarrow A\rightarrow E\stackrel{{\scriptstyle\eta}}{{\rightarrow}}G\rightarrow 1$ (see Definition 13), we define a homomorphism $(G,G)\to E$ by mapping the generators $\langle x,y\rangle$, $(x^{\prime},y^{\prime})$, and $(z)$ to $\overline{x},\overline{y}]$, $\overline{x}^{\prime},\overline{y}^{\prime}$}, and $\overline{z}^{2}$, respectively, where $\eta(\overline{x})=x$, $\eta(\overline{y})=y$, $\eta(\overline{x}^{\prime})=x^{\prime}$, $\eta(\overline{y}^{\prime})=y^{\prime}$, and $\eta(\overline{z})=z$. Because the extension is square-central, this homomorphism is well defined. Moreover, its restriction to $K^{\prime}$ has image contained in $S(E)\cap A$, and it sends $N^{\prime}$ to $1$, inducing a surjective homomorphism $\mathcal{M}(G;\mathbb{Z}_{2})\to S(E)\cap A$. Consequently, we obtain the exact sequence
@@ 4 The unoriented Bogomolov multiplier
For a finite group $G$, the unoriented Bogomolov multiplier, denoted $B_{0}(G;\mathbb{Z}_{2})$, is the quotient of the unoriented Schur multiplier $\mathcal{M}(G;\mathbb{Z}_{2})$ by the subgroup generated by the $G$-cobordisms over the torus $\langle x,y\rangle$, the Klein bottle $(x^{\prime},y^{\prime})$, and the projective space $(z)$, subject to $[x,y]=1$, $\{x^{\prime},y^{\prime}\}=1$, and $z^{2}=1$.
@@
The previous discussion implies that when a group $G$ has a Schur multiplier which is an elementary abelian $2$-group, then $B_{0}(G)=B_{0}(G;\mathbb{Z}_{2})$. This property holds for extraspecial $2$-groups of order $2^{2n+1}$ with $n>1$, where the Schur multiplier is an elementary abelian $2$-group of order $2^{2n^{2}-n-1}$; see [Kar87, p. 138].
@@
For finite abelian, dihedral, symmetric groups, and alternating groups the unoriented Bogomolov multiplier is trivial.
@@
Using [GAP], one can verify that the first integral homology are the image of two subgroups isomorphic to $\mathbb{Z}_{4}$ and one $\mathbb{Z}_{8}$. Moreover, the image of the second integral homology of every proper subgroup is trivial. Consequently, the class (51) remains nontrivial in the unoriented Bogomolov multiplier $B_{0}(\mathbb{Z}_{8}\rtimes Q_{8};\mathbb{Z}_{2})$. The reader may observe that, by applying the isomorphism (50) and that $B_{0}(G)$ is an elementary abelian 2-group, one immediately concludes that the unoriented Bogomolov multiplier is nontrivial.